OpenStudy (anonymous):

3h-5h+11=18 I know the answer: Negative seven over two. But I cannot seem to figure out how they got it. Medals will be rewarded. :) Thanks!

OpenStudy (akashdeepdeb):

How much is 3h - 5h? ===> -2h Subtract 11 from both sides -2h + 11 - 11 = 18 - 11 -2h = 18-11 = 7 h = -7/2 Understood? :)

OpenStudy (anonymous):

Ohh okay! So let me get this straight: If there are two numbers left that don't go into each other, for example: 2x=3, the answer would be 2 over 3?

OpenStudy (akashdeepdeb):

If 2x = 3 And we only want the value of x then divide both sides by 2 Leaving us 2x/2 which is x because that's what we want... D: So if 2x = 3 x = 3/2 If say 2x = 3 - 4x Add 4x to both sides to get rid of the RHS 4x 2x + 4x = 3 - 4x + 4x 6x = 3 x = 1/2 Understood? :)